Suppose IQ scores were obtained for 20 randomly selected sets of twins. The 20 pairs of measurements yield x overbarequals100.75​, y overbarequals102.2​, requals0.880​, ​P-valueequals​0.000, and ModifyingAbove y with caret equals negative 2.5 plus 1.04 x​, where x represents the IQ score of the twin born first.

Find the best predicted value of ModifyingAbove y with caret given that the twin born first has an IQ of 109​? Use a significance level of 0.05.
Answer #1

p value=0.000 < alpha=0.05

so, this regression model is useful.

y= -2.5 + 1.04x

so,  best predicted value is

y= -2.5 + 1.04*109=110.86 (answer)